helloo i need some help imm here u go
1 Solve the equation for the indicated variable

r-s=t;r



i super need to know how to do this THAnks
(:

Answer by stanbon(75887)

r-s=t; r = ?
------
Add s to both sides of the equation::
r-s+s = t+s
-----
Since -s+s = 0 you get:
r + 0 = t+s
---
r = t+s
===============
Cheers,
Stan H.
===============